# Distributional Vision Language Alignment by Cauchy-Schwarz Divergence

### v2

@@ Abstract
Multimodal alignment is crucial for various downstream tasks such as cross-modal generation and retrieval. Previous multimodal approaches like CLIP utilize InfoNCE to maximize mutual information, primarily aligning pairwise samples across modalities while overlooking distributional differences. In addition, InfoNCE has inherent conflict in terms of alignment and uniformity in multimodality, leading to suboptimal alignment with modality gaps. To overcome the limitations, we propose CS-Aligner, a novel framework that performs distributional vision-language alignment by integrating Cauchy-Schwarz (CS) divergence with mutual information. CS-Aligner captures both the global distribution information of each modality and the pairwise semantic relationships. We find that the CS divergence seamlessly addresses the InfoNCE’s alignment-uniformity conflict and serves complementary roles with InfoNCE, yielding tighter and more precise alignment. Moreover, by introducing distributional alignment, CS-Aligner enables incorporating additional information from unpaired data and token-level representations, enhancing flexible and fine-grained alignment in practice. Experiments on text-to-image generation and cross-modality retrieval tasks demonstrate the effectiveness of our method on vision-language alignment.
@@ 1 Introduction
Modality alignment aims to map the paired modal inputs (e.g., text and image) into a shared feature space, enabling success across diverse applications such as image-text retrieval [1, 2], text-to-image (T2I) generation [3, 4], and multimodal chatbots [5]. As a pioneering work in this field, CLIP [6] leverages InfoNCE loss (a.k.a. contrastive loss) to maximize the mutual information between paired text and image representations, effectively capturing pairwise and semantic relationships. Its versatility has made it a foundation for many multimodal tasks [3, 7].
Despite its success, CLIP and its variants [8, 9] exhibit a persistent modality gap, a misalignment between text and image representations in the shared latent space. As shown in Fig. 1, text and image embeddings often fail to align precisely and may remain separated from each other. This phenomenon has been widely observed [10, 11, 12] and is attributed to issues such as cone effects [11] or suboptimal latent space geometry [12]. Intriguingly, Liang et al. [11] observed that CLIP’s InfoNCE loss could inadvertently exacerbate the modality gap, since, as analyzed in Sec. 2, InfoNCE loss can be decomposed into alignment and uniformity components, which indeed conflict with each other during multimodal alignment.
@@
Several strategies have been proposed to mitigate the modality gap, such as projection modules with cosine similarity [10, 13, 1] and geodesic multimodal mixup [14]. UnCLIP-based models like DALL-E 2 [3] and Kandinsky [4] employ text-to-image prior modules (e.g., diffusion models) to map text embeddings to image feature space for alignment. A more recent alternative Eclipse [15] uses $\ell_{2}$ loss to train a prior adapter for text and image alignment. These works aim to transform representations across modalities for alignment. However, they explore alignment sample-wisely, heavily relying on pairwise data. Although sample-wise alignment effectively captures semantic information, it falls short in aligning entire data distributions. Similar to the InfoNCE in CLIP, the methods struggle to match the representation spaces across modalities, ultimately limiting the overall alignment. The reliance on carefully curated text-image pairs also limits the scalability and applicability to real-world scenarios with unpaired and noisy datasets [16, 17]. Moreover, the theoretical conflict of InfoNCE for multimodal alignment is still under exploration.
To address these challenges, we propose CS-Aligner, a novel distributional approach that incorporates Cauchy-Schwarz (CS) divergence [18] for modality alignment. As a symmetric measure, CS divergence robustly and efficiently estimates the distance between any representation distributions without parametric distributional assumptions, making it highly suitable for multimodal distribution alignment. Furthermore, we analyze the inherent conflict between alignment and uniformity in InfoNCE within the multimodal setting, and find that CS divergence can effectively mitigate the conflict and remains compatible with InfoNCE by estimating via a kernel density estimator (KDE) [19]. This enables CS-Aligner to align vision and language representations in both distributional and sample-wise levels, considering both the global modality and local semantic information, leading to more comprehensive, consistent, and tighter alignment as shown in Fig. 1  and Fig. 1.
@@ 2 InfoNCE is insufficient for alignment
Previous multimodal methods like CLIP [6] learn text and image representations in a shared space by maximizing lower bounds (e.g., InfoNCE [20]) of mutual information between modalities:
@@
Limitation1: Mutual information is insufficient for multimodal alignment. Although widely adopted, mutual information alone is insufficient for effective modality alignment [11]. The reason is that mutual information quantifies the statistical dependence between two random variables [21], ensuring correlation maximization between two random variables. However, it does not guarantee that the distributions $p(\mathbf{x})$ and $p(\mathbf{y})$ are statistically similar or close to each other in terms of their underlying distributions. In other words, the embedding distributions of two modalities can differ significantly or be far apart, yet exhibit strong dependence.
We illustrate this issue using a toy example in Fig. 2. Fig. 2a shows that despite strong dependence and high mutual information, the representation distributions of two representations or random variables can remain misaligned and be far from each other, resulting in a high divergence. This issue is also observed in the CLIP model pretrained with InfoNCE, where the vision and language representations exhibit a noticeable distributional gap, as shown in Fig. 1. This gap results in inconsistently aligned multimodal features, hindering the clear representation of shared semantics and disrupting effective mapping between modalities. Ultimately, this misalignment degrades performance in downstream tasks, including cross-modality generation. Ideally, the desired multimodal representations should be highly correlated with low distributional divergence, as depicted in Fig. 2b. Notably, although directly minimizing the divergence between distributions may reduce the distributional gap, it risks creating independent multimodal distributions without common semantic information (Fig. 2c). Therefore, maximizing mutual information and minimizing divergence complement each other to achieve effective multimodal representation alignment. Details are provided in Appendix B.
Limitation2: InfoNCE includes conflicting terms for multimodal alignment. In practice, mutual information is often optimized via minimizing the InfoNCE loss [20] which estimates $I(\mathbf{x};\mathbf{y})$ using paired image-text data $\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ and contains image-text and text-image alignment terms:
@@
As analyzed in [22], the InfoNCE loss can be decomposed as the sum of the alignment and uniformity terms i.e., $\mathcal{L}_{\text{InfoNCE}}\approx\mathcal{L}_{\text{align}}+\mathcal{L}_{\text{uniform}}$:
@@
[…]
where $t$ and $\alpha$ are hyperparameters. $p_{\text{pair}}$ denotes the image-text pairs distribution. Minimizing $\mathcal{L}_{\text{align}}$ encourages pairwise alignment. In unimodality, minimizing $\mathcal{L}_{\text{uniform}}$ promotes representations that are uniformly distributed on the unit hypersphere, a desirable property for representation learning [22]. However, in multimodal alignment, $\mathcal{L}_{\text{uniform}}$ may conflict with $\mathcal{L}_{\text{align}}$.
@@
$\mathcal{L}_{\text{uniform}}\approx-t\mathbb{E}_{\mathbf{x},\mathbf{y}\sim p(\mathbf{x},\mathbf{y})}\left[||\mathbf{x}-\mathbf{y}||_{2}^{2}\right]=-t\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p_{\text{pair}}+p_{\text{unpair}}}\left[||\mathbf{x}-\mathbf{y}||_{2}^{2}\right],$ (4)
@@
$\mathcal{L}_{\text{align}}+\mathcal{L}_{\text{uniform}}\approx\mathbb{E}_{(x,y)\sim p_{\text{pair}}}\left[||\mathbf{x}-\mathbf{y}||_{2}^{\alpha}\right]-t\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p_{\text{pair}}+p_{\text{unpair}}}\left[||\mathbf{x}-\mathbf{y}||_{2}^{2}\right].$ (5)
@@ 3.1 CS-Aligner: distributional multimodal alignment
To this end, we propose a distributional alignment framework, namely CS-Aligner, which leverages the Cauchy-Schwarz divergence ($D_{\text{CS}}$[23, 18], as illustrated in Fig. 3. The objective is:
@@
[…]
@@
CS divergence estimation. To estimate CS divergence, we introduce non-parametrical KDE [19], eliminating the need for explicit parametric assumptions about the underlying distributions. This provides significant flexibility in measuring distributional distance. Given i.i.d. samples $\{\mathbf{x}_{i}\}_{i=1}^{M}\sim p(\mathbf{x})$ and $\{\mathbf{y}_{i}\}_{i=1}^{N}\sim p(\mathbf{y})$, the empirical CS divergence estimator is given by [24]:
@@
When estimating the mutual information term $I(\mathbf{x},\mathbf{y})$ via InfoNCE (Eq. (2)), unlike other distribution divergences, CS divergence effectively addresses InfoNCE’s inherent alignment-uniformity conflict.
@@
[…]
@@
The connection between CS divergence and InfoNCE becomes evident when analyzing both terms from a cosine similarity perspective. For a characteristic kernel $\kappa(\mathbf{x},\mathbf{y})=\langle\phi(\mathbf{x}),\phi(\mathbf{y})\rangle_{\mathcal{H}}$, where $\phi$ maps samples to a Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}$, the mean embeddings are: $\boldsymbol{\mu}_{x}=\frac{1}{m}\sum_{i=1}^{m}\phi(\mathbf{x}_{i})\quad\text{and}\quad\boldsymbol{\mu}_{y}=\frac{1}{n}\sum_{i=1}^{n}\phi(\mathbf{y}_{i}),$ The CS divergence can then be expressed as:
$\hat{D}_{\mathrm{CS}}(p(\mathbf{x});p(\mathbf{y}))$ $=-2\log\left(\frac{\langle\boldsymbol{\mu}_{x},\boldsymbol{\mu}_{y}\rangle_{\mathcal{H}}}{\|\boldsymbol{\mu}_{x}\|_{\mathcal{H}}\|\boldsymbol{\mu}_{y}\|_{\mathcal{H}}}\right)=-2\log\text{sim}(\boldsymbol{\mu}_{x},\boldsymbol{\mu}_{y}),$ (11)
which evaluates the cosine similarity between distributions in RKHS. Similarly, InfoNCE evaluates cosine similarity between paired samples (Eq. (2)). This dual-level similarity assessment underscores the synergy between CS divergence and mutual information, offering a unified and robust framework for multimodal alignment.
@@ 3.2 Extend CS-Aligner to unpaired data
Benefits from the distributional alignment, we further propose extensions of CS-Aligner, which leverage additional information in unpaired data. While the mutual information estimation (InfoNCE) part requires pairwise data, the CS divergence estimator (Eq. (8)) can operate seamlessly on unpaired data without introducing additional computation. This unique capability enables CS-Aligner to extend beyond traditional pairwise multimodal alignment by incorporating additional distributional information from unpaired data or tokens. Below, we introduce two novel directions.
@@
Vision-language token alignment. We propose a novel intra-sample distribution alignment approach between vision and language tokens. Unlike CLIP-based models [6] which align only the “CLS" tokens of vision and text representations, the method considers all vision and text tokens for a more fine-grained alignment. Specifically, each vision feature $\mathbf{x}_{i}\in\mathbb{R}^{V\times D}$ is modeled as a token distribution $p(\mathbf{x}_{i})$ containing $V$ vision tokens, while each text feature $\mathbf{y}_{i}\in\mathbb{R}^{L\times D}$ is represented as a token distribution $p(\mathbf{y}_{i})$ with $L$ text tokens. $D$ denotes the feature dimension. We compute CS divergence between vision and text token distributions, and obtain internal token-wise alignment loss:
@@ 3.3 Parameter-efficient multimodal alignment
We demonstrate the effectiveness of our CS-Aligner by performing vision-language alignment in a parameter-efficient manner using pretrained vision and language models, such as CLIP and large language models (LLMs) [25]. To adapt these pretrained models, we employ two widely used frameworks: adapter [13] and LoRA [26]. The adapter and LoRA enable efficient alignment of the multimodal large-scale pretrained models, without requiring extensive computational resources. The whole framework is demonstrated in Fig. 3.
Adapter alignment. We add a lightweight transformer [27] on top of the pretrained model as an adapter. The adapter projects text embeddings or image embeddings into a shared space.
LoRA alignment. We insert trainable low-rank matrices into the pretrained weights of text encoder. It enables fine-grained adjustments to the representations, aligning them with the other modality.

### v3

@@ Abstract
Vision-language alignment is crucial for various downstream tasks such as cross-modal generation and retrieval. Previous multimodal approaches like CLIP utilize InfoNCE to maximize mutual information, primarily aligning pairwise samples across modalities while overlooking distributional differences. In addition, InfoNCE has inherent conflict in terms of alignment and uniformity in multimodality, leading to suboptimal alignment with modality gaps. To overcome the limitations, we propose CS-Aligner, a novel framework that performs distributional vision-language alignment by integrating Cauchy-Schwarz (CS) divergence with mutual information. CS-Aligner captures both the global distribution information of each modality and the pairwise semantic relationships. We find that the CS divergence seamlessly addresses the InfoNCE’s alignment-uniformity conflict and serves complementary roles with InfoNCE, yielding tighter and more precise alignment. Moreover, by introducing distributional alignment, CS-Aligner enables incorporating additional information from unpaired data and token-level representations, enhancing flexible and fine-grained alignment in practice. Experiments on text-to-image generation and cross-modality retrieval tasks demonstrate the effectiveness of our method on vision-language alignment.
@@ 1 Introduction
Vision-language alignment aims to map the paired text and image inputs into a shared feature space, enabling success across diverse applications such as image-text retrieval (Huang et al., 2024; Koukounas et al., 2024) and text-to-image (T2I) generation (Ramesh et al., 2022; Razzhigaev et al., 2023). As a pioneering work in this field, CLIP (Radford et al., 2021) leverages InfoNCE loss (a.k.a. contrastive loss) to maximize the mutual information between paired text and image representations, effectively capturing pairwise and semantic relationships. Its versatility has made it a foundation for many multimodal tasks (Ramesh et al., 2022; Mokady et al., 2021).
Despite its success, CLIP and its variants (Zhai et al., 2023; Sun et al., 2023) exhibit a persistent modality gap, a misalignment between text and image representations in the shared latent space. As shown in Fig. 1, text and image embeddings often fail to align precisely and may remain separated from each other. This phenomenon has been widely observed (Zhou et al., 2023; Liang et al., 2022; Shi et al., 2023) and is attributed to issues such as cone effects (Liang et al., 2022) or suboptimal latent space geometry (Shi et al., 2023). Intriguingly, Liang et al. (Liang et al., 2022) observed that CLIP’s InfoNCE loss could inadvertently exacerbate the modality gap, since, as analyzed in Sec. 2, InfoNCE loss can be decomposed into alignment and uniformity components, which indeed conflict with each other during vision-language alignment.
@@
Several strategies have been proposed to mitigate the modality gap, such as projection modules with cosine similarity (Zhou et al., 2023; Gao et al., 2024; Huang et al., 2024) and geodesic multimodal mixup (Oh et al., 2024). UnCLIP-based models like DALL-E 2 (Ramesh et al., 2022) employ text-to-image prior modules (e.g., diffusion models) to map text embeddings to image feature space for alignment. A more recent alternative Eclipse (Patel et al., 2024) uses $\ell_{2}$ loss to train a prior adapter for text and image alignment. These works aim to transform representations across modalities for alignment. However, they explore alignment sample-wisely, heavily relying on pairwise data. Although sample-wise alignment effectively captures semantic information, it falls short in aligning entire data distributions. Similar to the InfoNCE in CLIP, the methods struggle to match the representation spaces across modalities, ultimately limiting the overall alignment. The reliance on carefully curated text-image pairs also limits the scalability and applicability to real-world scenarios with unpaired and noisy datasets (Lin et al., 2014; Li et al., 2023b). Moreover, the theoretical conflict of InfoNCE for vision-language alignment is still under exploration.
To address these challenges, we propose CS-Aligner, a novel distributional approach that incorporates Cauchy-Schwarz (CS) divergence (Principe et al., 2000b) for vision-language alignment. As a symmetric measure, CS divergence robustly and efficiently estimates the distance between any representation distributions without parametric distributional assumptions, making it highly suitable for multimodal distribution alignment. Furthermore, we analyze the alignment–uniformity conflict of InfoNCE in multimodal settings and show that CS divergence effectively mitigates it while remaining compatible with InfoNCE via kernel density estimation (KDE) (Parzen, 1962). This enables CS-Aligner to align vision–language representations at distributional and sample-wise levels, capturing global modality and local semantics, yielding more comprehensive, consistent, and tighter alignment as shown in Figs. 1 and 1.
@@ 2 InfoNCE is insufficient for alignment
Previous multimodal methods (for vision-language) like CLIP (Radford et al., 2021) learn text and image representations in a shared space by maximizing lower bounds (e.g., InfoNCE (Oord et al., 2018)) of mutual information between modalities:
@@
Limitation1: Mutual information is insufficient for multimodal alignment. Although widely adopted, mutual information alone is insufficient for effective modality alignment (Liang et al., 2022). The reason is that mutual information quantifies the statistical dependence between two random variables (Cover, 1999), ensuring correlation maximization between two random variables. However, it does not guarantee that the distributions $p(\mathbf{x})$ and $p(\mathbf{y})$ are statistically similar or close to each other in terms of their underlying distributions. In other words, the embedding distributions of two modalities can differ significantly or be far apart, yet exhibit strong dependence.
We illustrate this issue using a toy example in Fig. 2. Fig. 2a shows that despite strong dependence and high mutual information, the representation distributions of two representations or random variables can remain misaligned and be far from each other, resulting in a high divergence. This issue is also observed in the CLIP model pretrained with InfoNCE, where the vision and language representations exhibit a noticeable distributional gap, as shown in Fig. 1. This gap results in inconsistently aligned multimodal features, hindering the clear representation of shared semantics and disrupting effective mapping between modalities. Ultimately, this misalignment degrades performance in downstream tasks, including cross-modality generation. Ideally, the desired multimodal representations should be highly correlated with low distributional divergence, as depicted in Fig. 2b. Notably, although directly minimizing the divergence between distributions may reduce the distributional gap, it risks creating independent multimodal distributions without common semantic information (Fig. 2c). Therefore, maximizing mutual information and minimizing divergence complement each other to achieve effective multimodal representation alignment. Details are provided in Appendix A.
Limitation2: InfoNCE includes conflicting terms for multimodal alignment. In practice, mutual information is often optimized via the InfoNCE loss (Oord et al., 2018) which estimates $I(\mathbf{x};\mathbf{y})$ using paired image-text data $\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ and contains image-text and text-image alignment terms:
@@
As analyzed in Wang and Isola (2020), the InfoNCE loss can be decomposed as the sum of the alignment ($\mathcal{L}_{\text{align}}$) and uniformity ($\mathcal{L}_{\text{uniform}}$) terms i.e., $\mathcal{L}_{\text{InfoNCE}}\approx\mathcal{L}_{\text{align}}+\mathcal{L}_{\text{uniform}}$:
@@
where $t$ and $\alpha$ are hyperparameters. $p_{\text{pair}}$ denotes the image-text pairs distribution. Minimizing $\mathcal{L}_{\text{align}}$ encourages pairwise alignment. In unimodality, minimizing $\mathcal{L}_{\text{uniform}}$ promotes representations that are uniformly distributed on the unit hypersphere, a desirable property for representation learning (Wang and Isola, 2020). However, in multimodal alignment, $\mathcal{L}_{\text{uniform}}$ may conflict with $\mathcal{L}_{\text{align}}$.
@@
[…]
@@
[…]
@@ 3.1 CS-Aligner: distributional multimodal alignment
To this end, we propose a distributional alignment framework, namely CS-Aligner, which leverages the CS divergence ($D_{\text{CS}}$), as illustrated in Fig. 3. The objective is:
@@
$D_{\text{CS}}(p;q)=-\log\Big({\Big(\int p(\mathbf{\omega})q(\mathbf{\omega})d\mathbf{\omega}\Big)^{2}}/\ \Big({\int p(\mathbf{\omega})^{2}d\mathbf{\omega}\int q(\mathbf{\omega})^{2}d\mathbf{\omega}}\Big)\Big),$ (7)
@@
CS divergence estimation. To estimate CS divergence, we introduce non-parametrical KDE. The non-parametric KDE means that it does not assume any specific parametric form for the underlying distribution. This eliminates the need for explicit parametric assumptions about the underlying distributions. This provides significant flexibility in measuring distributional distance. Given i.i.d. samples $\{\mathbf{x}_{i}\}_{i=1}^{M}\sim p(\mathbf{x})$ and $\{\mathbf{y}_{i}\}_{i=1}^{N}\sim p(\mathbf{y})$, the empirical CS divergence estimator is given by Jenssen et al. (2006):
@@
When estimating the mutual information $I(\mathbf{x},\mathbf{y})$ via InfoNCE (Eq. (2)), unlike other distribution divergences, CS divergence effectively addresses InfoNCE’s inherent alignment-uniformity conflict.
@@
$\;+\lambda\Bigl(\log\,\mathbb{E}_{\mathbf{x},\mathbf{x}^{\prime}\sim p(\mathbf{x})}\bigl[\kappa_{t}(\mathbf{x},\mathbf{x}^{\prime})\bigr]+\log\,\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim p(\mathbf{y})}\bigl[\kappa_{t}(\mathbf{y},\mathbf{y^{\prime}})\bigr]-2\,\log\,\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x}),\,\mathbf{y}\sim p(\mathbf{y})}\bigl[\kappa_{t}(\mathbf{x},\mathbf{y})\bigr]\Bigr).$
@@
The connection between CS divergence and InfoNCE becomes evident when analyzing both terms from a cosine similarity perspective. For a characteristic kernel $\kappa(\mathbf{x},\mathbf{y})=\langle\phi(\mathbf{x}),\phi(\mathbf{y})\rangle_{\mathcal{H}}$, where $\phi$ maps samples to a Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}$, the mean embeddings are: $\bm{\mu}_{x}=\frac{1}{m}\sum_{i=1}^{m}\phi(\mathbf{x}_{i})\quad\text{and}\quad\bm{\mu}_{y}=\frac{1}{n}\sum_{i=1}^{n}\phi(\mathbf{y}_{i}),$ The CS divergence can then be expressed in a form that evaluates the cosine similarity between distributions in RKHS:
$\hat{D}_{\mathrm{CS}}(p(\mathbf{x});p(\mathbf{y}))$ $=-2\log\left(\frac{\langle\bm{\mu}_{x},\bm{\mu}_{y}\rangle_{\mathcal{H}}}{\|\bm{\mu}_{x}\|_{\mathcal{H}}\|\bm{\mu}_{y}\|_{\mathcal{H}}}\right)=-2\log\text{sim}(\bm{\mu}_{x},\bm{\mu}_{y}),$ (11)
Similarly, InfoNCE evaluates cosine similarity between paired samples (Eq. (2)). This dual-level similarity assessment underscores the synergy between CS divergence and mutual information, offering a unified and robust framework for multimodal alignment.
@@ 3.2 Extend CS-Aligner to unpaired data
Benefiting from the distributional alignment, we further propose extensions of CS-Aligner, which leverage additional information in unpaired data. While the mutual information estimation (InfoNCE) part requires pairwise data, the CS divergence estimator (Eq. (8)) can operate seamlessly on unpaired data without introducing additional computation. This unique capability enables CS-Aligner to extend beyond traditional pairwise multimodal alignment by incorporating additional distributional information from unpaired data or tokens. Below, we introduce two novel directions.
@@
Vision-language token alignment. We propose a novel intra-sample distribution alignment approach between vision and language tokens. Unlike CLIP-based models (Radford et al., 2021) aligning only the “CLS” tokens of vision and text, our method aligns all tokens for finer-grained alignment. Specifically, each vision feature $\mathbf{x}_{i}\in\mathbb{R}^{V\times D}$ is modeled as a token distribution $p(\mathbf{x}_{i})$ containing $V$ vision tokens, while each text feature $\mathbf{y}_{i}\in\mathbb{R}^{L\times D}$ is represented as a token distribution $p(\mathbf{y}_{i})$ with $L$ text tokens. $D$ denotes the feature dimension. We compute CS divergence between vision and text token distributions, and obtain an internal token-wise alignment loss:
@@ 3.3 Parameter-efficient multimodal alignment
We demonstrate the effectiveness of our CS-Aligner by performing vision-language alignment in a parameter-efficient manner using pretrained vision and language models, such as CLIP and large language models (LLMs) (Dubey et al., 2024). To adapt these pretrained models, we employ two widely used frameworks: adapter (Gao et al., 2024) and LoRA (Hu et al., 2021). The adapter and LoRA enable efficient alignment of the multimodal large-scale pretrained models, without requiring extensive computational resources. The whole framework is demonstrated in Fig. 3.
Adapter & LoRA alignment. We add a lightweight transformer (Vaswani, 2017) on top of the pretrained model as an adapter that projects text or image embeddings into a shared space; optionally, we can insert trainable low-rank (LoRA) matrices into the text encoder’s weights to enable fine-grained adjustments, aligning the representations with the other modality.